OpenStudy (anonymous):

Simplify 9 and 1/3 times 8 and 1/1^3

sammixboo (sammixboo):

\(9\frac{1}{3}\times 8\frac{1}{1^3}\) ?

OpenStudy (anonymous):

\[9\frac{ 1 }{ 3 }\times8\frac{ 1 }{ 1^{3} }=9\frac{ 1 }{ 3 }\times8\frac{ 1 }{ 1 }=9\frac{ 1 }{ 3 }\times9=84\]